Hi
Do i cross multiply first or do i find a common demonator? i tried both but i still got lost.
Can someone help me with this problem? x+1 /3+x+2 /7=5
I would really appreciate it if someone would help me solve this
Thank You

Answer by stanbon(75887)

(x+1)/3+(x+2)/7 = 5
Multiply thru by 21 to get:
7(x+1) + 3(x+2) = 105
7x+7+3x+6 = 105
10x+13=105
10x = 92
x=46/5= 9.2
Cheers,
Stan H.